OpenStudy (anonymous):

can anyone explain some trigonometry for me? mainly how to do this: a)Use a sum or difference formula to calculate the exact value of sin( pi/12 ). b) Use part a) to determine the exact value of cos(5pi/12 ).

OpenStudy (anonymous):

pi/12=pi/3-pi/4 now use it

OpenStudy (anonymous):

i must be tired...

OpenStudy (anonymous):

cos(5pi/2)=cos(pi/2-pi/12)=sin(pi/12)

OpenStudy (anonymous):

sin pi/3 cos pi/4 - cos pi/3 sin pi/4

OpenStudy (anonymous):

yes